### v1

# WIMP/FIMP dark matter and primordial black holes with memory burden effect

Abstract
The lifetime of primordial black holes (PBHs), which were formed in the early universe, can be extended by the memory burden effect. Light PBHs may be in existence today and become a candidate for dark matter (DM). We assume that DM is made of thermally produced weakly interacting massive particles (WIMPs), WIMPs produced via Hawking radiation of PBHs, and survived PBHs via the memory burden effect. In addition, feebly interacting massive particles (FIMPs) are considered as an alternative to WIMPs. Various similar studies have been conducted either with or without considering the memory burden effect. In this study, we simultaneously account for thermally produced WIMPs or FIMPs and the memory burden effect of PBHs to explain the relic abundance of DM. We show that this abundance is highly sensitive to the memory burden effect in PBHs.
1 Introduction
The existence and production mechanism of dark matter (DM) have been a longstanding mystery in cosmology and particle physics [1]. The traditional and most studied thermal DM production mechanism is freeze-out (FO) [2]. In this mechanism, DM particles $\chi$ reach thermal equilibrium with thermal bath particles $f$ in the early universe through annihilation and pair production $\chi\chi\leftrightarrow\bar{f}f$. Subsequently, the temperature of bath particles drops below the mass of DM particles and $\Gamma\lesssim H$ is satisfied, where $\Gamma$ is the reaction rate of annihilation $\chi\chi\rightarrow\bar{f}f$ and $H$ denotes the Hubble parameter. Then, the comoving number density of DM particles remains constant. As a typical weak-scale cross-section leads to the observed relic abundance of DM, the FO scenario is usually discussed in the context of weakly interacting massive particles (WIMPs). Although WIMPs remain a leading candidate explanation for DM, they elude detection in numerous experiments. Thus, other candidates have been proposed, such as feebly interacting massive particles (FIMPs) [3, 4] based on the opposite idea of WIMPs. If the coupling of DM particles $\chi$ to bath particles is very weak, DM particles never reach thermal equilibrium with the thermal bath. Such a DM particle is called a FIMP. Assuming that the initial density of FIMPs is negligible, the feeble interaction with the thermal bath leads to FIMP production during the expansion of the universe. This DM production scheme is called the freeze-in (FI) mechanism and is applicable to different scenarios such as decay $f_{1}\rightarrow f_{2}\chi$, scattering $f_{1}f_{2}\rightarrow f_{3}\chi$, and pair production $f_{1}f_{2}\rightarrow\chi\chi$.
Primordial black holes (PBHs), which were formed in the early universe, are also feasible candidates for explaining DM [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. PBHs emit particles through Hawking radiation [17], which reduces their mass. Therefore, PBHs have a finite lifetime. However, if this lifetime is longer than the age of the universe, PBHs may exist until the present day without being completely evaporated.
PBHs that formed with a mass of less than approximately $10^{15}$ g were believed not to survive to the present day. Nevertheless, it has recently been discovered that the lifetime of PBHs can be extended by the memory burden effect [18, 19, 20, 21, 22]. The quantum corrections for PBHs due to the memory burden effect are being extensively studied to explain the relic abundance of PBH [23, 26, 24, 25], axion DM [27], gravitational waves [28, 29, 26], baryon asymmetry [32, 30, 31], cosmic neutrinos [33, 34, 35], compact stars [36], and phenomena beyond the Standard Model of particle physics [37].
We assume that DM is made of both WIMPs (or FIMPs) and survived PBHs. As Hawking radiation is induced by gravity, PBHs evaporate into all particle species. WIMPs (FIMPs) are also produced by Hawking radiation. Thus, we should treat the following three DM components:
  • Thermally produced WIMPs (or FIMPs),
  • WIMPs (or FIMPs) produced from Hawking radiation of PBHs, and
  • Survived PBHs.
Similar studies have neglected the memory burden effect [39, 38, 40, 41, 42, 43, 44, 45, 46, 47]. Considering the memory burden effect, a study attempted to explain the relic abundance of DM via Hawking radiation of PBHs and survived PBHs [23]. However, this study did not account for DM related to thermally produced particles.
We simultaneously consider thermally produced WIMPs/FIMPs and the memory burden effect of PBHs to explain the relic abundance of DM. If the effect of thermal production of WIMPs or FIMPs is much smaller than that of the particle production by PBHs via Hawking radiation, we can omit the contribution of the thermally produced WIMPs or FIMPs. In this case, the results are the same as those reported in [23]. Therefore, we focus on the case that the effect of thermal production of WIMPs or FIMPs is much larger than the effect of the particle production by PBHs via Hawking radiation.
The remainder of this paper is organized as follows. Section 2 provides a brief review of PBHs with the memory burden effect. In Sec. 3, we assess the combined effects of PBHs and thermally produced WIMPs or FIMPs on the relic abundance of DM. Finally, Sec. 4 provides a summary of the study and findings.
2 PBH with memory burden effect
2.1 Time evolution
Neglecting the memory burden effect, the time evolution of PBH mass $M_{\rm BH}$ is given by
$\frac{dM_{\rm BH}}{dt}=-\epsilon\frac{M_{\rm P}^{4}}{M_{\rm BH}^{2}},$ (1)
where $M_{\rm P}\simeq 2.4\times 10^{18}$ GeV is the reduced Planck mass and
$\epsilon=\frac{27}{4}\frac{\pi g_{\ast}(T_{\rm BH})}{480},$ (2)
with $g_{\ast}(T_{\rm BH})$ being the effective number of degrees of freedom. The Hawking temperature, $T_{\rm BH}$, is defined as
$T_{\rm BH}=\frac{M_{\rm P}^{2}}{M_{\rm BH}}.$ (3)
After integrating Eq. (1) from the time of PBH formation, $t_{\rm in}$, to time $t$, the PBH mass at time $t$ is calculated as
$M_{\rm BH}=M_{\rm in}\left[1-\Gamma_{\rm BH}^{0}(t-t_{\rm in})\right]^{1/3},$ (4)
where $M_{\rm in}$ is the initial PBH mass. The decay width associated with PBH evaporation is estimated to be
$\Gamma_{\rm BK}^{0}=\frac{3\epsilon M_{\rm P}^{4}}{M_{\rm in}^{3}}.$ (5)
Assuming PBH formation in a radiation-dominated era, PBH lifetime $\tau^{0}$ is obtained as
$\tau^{0}=t_{\rm ev}^{0}-t_{\rm in}=\frac{1}{\Gamma_{\rm BK}^{0}}\simeq 2.4\times 10^{-28}\left(\frac{M_{\rm in}}{1{\rm g}}\right)^{3}\ {\rm s},$ (6)
by solving $M_{\rm BH}(t_{\rm ev}^{0})=0$, where $t_{\rm ev}^{0}$ is the evaporation time.
Considering the memory burden effect, the time evolution of the PBH mass is modified to become [23]
$\frac{dM_{\rm BH}}{dt}=-\frac{\epsilon}{[S(M_{\rm BH})]^{k}}\frac{M_{\rm P}^{4}}{M_{\rm BH}^{2}},$ (7)
where
$S=\frac{1}{2}\left(\frac{M_{\rm BH}}{M_{\rm P}}\right)^{2}=\frac{1}{2}\left(\frac{M_{\rm P}}{T_{\rm BH}}\right)^{2}$ (8)
is the black hole entropy. Parameter $k$ characterizes the efficiency of the back reaction effect, with $k=0$ representing the case of Hawking radiation without the memory burden effect.
We assume that the semiclassical era is valid until the PBH mass reaches
$M_{\rm BH}=qM_{\rm in},$ (9)
with $10<q<$ 1. From Eq. (4), we define the final time for the semiclassical era as
$t_{q}=t_{\rm ev}^{0}(1-q^{3}).$ (10)
After integrating Eq. (7) from time $t_{q}$ to $t$, we obtain
$M_{\rm BH}=qM_{\rm in}\left[1-\Gamma_{\rm BH}^{k}(t-t_{q})\right]^{1/(3+2k)},$ (11)
where
$\Gamma_{\rm BH}^{k}=2^{k}(3+2k)\epsilon M_{\rm P}\left(\frac{M_{P}}{qM_{\rm in}}\right)^{3+2k}$ (12)
for $k\neq 0$ and $10<q<$ 1. We estimate the evaporation time by solving $M_{\rm BH}(t_{\rm ev}^{k})=0$ as follows [20, 23, 32]:
$t_{\rm ev}^{k}=t_{q}+\frac{1}{\Gamma_{\rm BK}^{k}}\simeq(1-q^{3})\frac{M_{\rm in}^{3}}{3\epsilon M_{\rm P}^{4}}+\frac{1}{2^{k}(3+2k)\epsilon M_{\rm P}}\left(\frac{qM_{\rm in}}{M_{P}}\right)^{3+2k}.$ (13)
Approximation $t_{\rm ev}^{k}\simeq 1/\Gamma_{\rm BH}^{k}$ is adequate in some cases.
2.2 Initial density
The initial density of a PBH can be described by the following ratio:
$\beta=\frac{\rho_{\rm BH}^{\rm in}}{\rho_{\rm r}^{\rm in}},$ (14)
where $\rho_{\rm BH}^{\rm in}$ and $\rho_{\rm r}^{\rm in}=\pi^{2}g_{\ast}(T_{\rm in})T_{\rm in}^{4}/30$ are the initial energy densities of the PBH and radiation, respectively. We assume that the PBH formed during a radiation-dominated era in the universe. Hence, the radiation energy density varies as follows:
$\rho_{\rm r}(a)=\rho_{\rm r}^{\rm in}\left(\frac{a_{\rm in}}{a}\right)^{4}.$ (15)
As a PBH behaves as matter, its energy density evolves as follows:
$\rho_{\rm BH}(a)=\rho_{\rm BH}^{\rm in}\left(\frac{a_{\rm in}}{a}\right)^{3}\times\begin{cases}1&{\rm for}\ a<a_{\rm q}\\ q&{\rm for}\ a>a_{\rm q}\\ \end{cases},$ (16)
where $a$ is a scale factor, and instant mass reduction is assumed as $M_{\rm in}\rightarrow qM_{\rm in}$ at $a=a_{\rm q}$. From $\rho_{\rm r}=\rho_{\rm BH}$, the radiation–PBH equivalent may occur at $a_{\rm BH}=a_{\rm in}/(q\beta)$. A PBH-dominant era is possible if $a_{\rm BH}<a_{\rm ev}$, which corresponds to
$\beta>\frac{a_{\rm in}}{qa_{\rm ev}}=\beta_{\rm c},$ (17)
where
$\beta_{\rm c}=\frac{1}{q^{5/2+k}}\left(\frac{M_{\rm P}}{M_{\rm in}}\right)^{1+k}\sqrt{\frac{(3+2k)2^{k}\epsilon}{8\pi\gamma}},$ (18)
with $\gamma\simeq 0.2$. We assume that the initial PBH mass is related to the horizon size at PBH formation as follows:
$M_{\rm in}=\frac{4}{3}\pi\gamma\rho_{\rm r}^{\rm in}H_{\rm in}^{-3}.$ (19)
In addition, we use relations $H=1/(2t)$, $a\propto t^{1/2}$, and $H_{\rm in}^{2}=\rho_{\rm r}^{\rm in}/(3M_{\rm P}^{2})$ for a radiation-dominated universe, and $t_{\rm ev}^{k}\simeq 1/\Gamma_{\rm BK}^{k}$ using Eq. (12).
As addressed in [23], $\beta<\beta_{c}$ establishes a case of interest because we assume that PBHs evaporate completely after the present time due to the memory burden effect and survived PBHs until the present day are candidates of DM (no early matter-dominated era is considered). Hereafter, we require relation $\beta<\beta_{c}$ to hold.
2.3 Survived PBH density
We assume that PBHs evaporate completely after the present time due to the memory burden effect. Thus, the density of survived PBHs is obtained as follows [48]:
$\Omega_{\rm PBH}h^{2}=1.6\times 10^{8}\frac{g_{0}}{g_{\rm q}}\frac{\rho_{\rm BH}(a_{\rm q})}{T^{3}(a_{\rm q})}\ {\rm GeV^{-1}}.$ (20)
Using relation
$\frac{\rho_{\rm BH}(a_{\rm q})}{T^{3}(a_{\rm q})}=\frac{q\rho_{\rm BH}^{\rm in}}{T_{\rm in}^{3}}=\frac{q\beta\rho_{r}^{\rm in}}{T_{\rm in}^{3}}$ (21)
and Eq. (19), we obtain (see Eq. (3.37) in [23])
$\Omega_{\rm PBH}h^{2}=4.2\times 10^{26}q\beta\left(\frac{M_{\rm P}}{M_{\rm in}}\right)^{1/2}.$ (22)
2.4 Particle production
The number of particle species, $i$, from PBH evaporation is controlled by the relation between particle mass $m_{i}$ and the PBH formation temperature given by
$T_{\rm in}=T_{\rm BH}(t_{\rm in})=\frac{M_{\rm P}^{2}}{M_{\rm in}}.$ (23)
If $m_{i}<T_{\rm in}$, particle emission from a PBH occurs from $t_{\rm in}$ to $t_{\rm ev}$. If $m_{i}>T_{\rm in}$, particle emission starts from time $t_{i}$, where $m_{i}=T_{\rm BH}(t_{i})$. The relic abundance of emitted particle DM is given by Eq. (3.33) in [23]:
$\Omega_{\rm ev}h^{2}=1.8\times 10^{8}N_{\chi}\beta\left(\frac{M_{\rm P}}{M_{\rm in}}\right)^{3/2}\frac{m_{\chi}}{\rm GeV},$ (24)
with
$N_{\chi}=\frac{15\xi g_{\chi}\zeta(3)}{g_{\ast}(T_{\rm BH})\pi^{4}}\times\begin{cases}(1-q^{2})\left(\frac{M_{\rm in}}{M_{\rm P}}\right)^{2},&{\rm for}\ m_{\chi}<T_{\rm in}\\ \left(\frac{M_{\rm P}}{m_{\chi}}\right)^{2}-\left(\frac{qM_{\rm in}}{M_{\rm P}}\right)^{2},&{\rm for}\ m_{\chi}>T_{\rm in}\\ \end{cases},$ (25)
where $\xi=1$ for bosons and $\xi=3/4$ for fermions.
3 WIMP/FIMP DM and PBHs
3.1 WIMPs and PBHs
First, we consider DM made of WIMPs and PBHs. FO of WIMPs occurs at [2]
$x_{\rm FO}=\frac{m_{\chi}}{T_{\rm FO}}\simeq 20,$ (26)
where $T_{\rm FO}$ is the FO temperature. We assume that FO occurs during a radiation-dominated era. The Hubble parameter is related to the radiation temperature and time by relation $H=(\pi^{2}g_{\ast}/30)T^{2}/(3M_{\rm P})=1/(2t)$. Thus, the FO time is obtained as follows:
$t_{\rm FO}=\frac{1}{2}\sqrt{\frac{90}{\pi^{2}g_{\ast}}}\frac{M_{\rm P}}{T_{\rm FO}^{2}}.$ (27)
The relic abundance of WIMPs $\chi$ via FO is calculated as [38]
$\Omega_{\rm FO}h^{2}=0.76\frac{x_{\rm FO}}{g_{\ast}^{1/2}}\frac{s_{0}}{M_{\rm P}\rho_{\rm c}\langle\sigma v\rangle}h^{2},$ (28)
where $\langle\sigma v\rangle$ is the thermally averaged cross-section for $\chi\chi\rightarrow\bar{f}f$ and $s_{0}=2.89\times 10^{3}\ {\rm cm^{-3}}$ and $\rho_{\rm c}=3H_{0}^{2}/(8\pi G)=1.05h^{2}\times 10^{-5}\ {\rm GeV\ cm^{-3}}$ are the entropy and critical densities in the present day, respectively. Considering the so-called WIMP miracle, we parameterize
$\langle\sigma v\rangle=\frac{\alpha_{\rm WIMP}^{2}}{m_{\chi}^{2}},$ (29)
where $\alpha_{\rm WIMP}^{2}$ is the effective coupling for $\chi\chi\leftrightarrow\bar{f}f$. The relic abundance of WIMPs is given by
$\Omega_{\rm FO}h^{2}=1.7\times 10^{-10}\frac{1}{\alpha_{\rm WIMP}^{2}}\left(\frac{m_{\chi}}{\rm GeV}\right)^{2}.$ (30)
If the FO mechanism alone allows to produce DM, observed relic abundance $\Omega_{\rm DM}h^{2}=0.12$ [49] is consistent with $\alpha_{\rm WIMP}\simeq 0.026$ and $m_{\chi}\simeq 1$ TeV.
If the PBH emission of WIMP $\chi$ precedes FO of $\chi$, the WIMPs emitted by Hawking radiation may reach chemical equilibrium with bath particles due to the massive thermal production of WIMPs. In this case, WIMPs from PBH evaporation provide no additional contribution to the relic abundance of thermally produced WIMPs. In contrast, if PBHs evaporate after WIMP FO, WIMPs emitted from PBHs may neither thermalize with the bath particles nor annihilate with each other. In this case, the WIMPs produced from PBHs may contribute to the relic abundance of DM [38, 40].
We assume that the DM particle is a Majorana fermion (e.g., $\xi=3/4$ and $g_{\chi}=1$). In addition, we omit the negligible contribution of $g_{\chi}=1$ to the effective number of degrees of freedom and use $g_{*}=106.75$. The present-day density of DM can be obtained as follows:
$\Omega_{\rm DM}h^{2}$ =$=\begin{cases}\Omega_{\rm FO}h^{2}+\Omega_{\rm PBH}h^{2}&{\rm for}\ t_{\rm ev}^{k}<t_{\rm FO}\\ \Omega_{\rm FO}h^{2}+\Omega_{\rm PBH}h^{2}+\Omega_{\rm ev}h^{2}&{\rm for}\ t_{\rm FO}<t_{\rm ev}^{k}\\ \end{cases},$
$=1.7\times 10^{-10}\frac{1}{\alpha_{\rm WIMP}^{2}}\left(\frac{m_{\chi}}{\rm GeV}\right)^{2}+4.2\times 10^{26}q\beta\left(\frac{M_{\rm P}}{M_{\rm in}}\right)^{1/2}$
+$\quad+\begin{cases}0&{\rm for}\ t_{\rm ev}^{k}<t_{\rm FO}\\ 2.3\times 10^{5}C\beta\frac{m_{\chi}}{\rm GeV}&{\rm for}\ t_{\rm FO}<t_{\rm ev}^{k}\\ \end{cases},$ (31)
where the first–third terms are due to FO, survived PBHs, and PBH evaporation, respectively. Coefficient $C$ is defined as
$C=\begin{cases}(1-q^{2})\left(\frac{M_{\rm in}}{M_{\rm P}}\right)^{1/2},&{\rm for}\ m_{\chi}<T_{\rm in}\\ \left(\frac{M_{\rm P}}{M_{\rm in}}\right)^{3/2}\left[\left(\frac{M_{\rm P}}{m_{\chi}}\right)^{2}-\left(\frac{qM_{\rm in}}{M_{\rm P}}\right)^{2}\right],&{\rm for}\ m_{\chi}>T_{\rm in}\\ \end{cases}$ (32)
where $T_{\rm in}$ is given by Eq. (23).
We study the case of $\Omega_{\rm FO}h^{2}\gg\Omega_{\rm ev}h^{2}$. In the following numerical calculations, we assume $k=1$ as the simplest case. As a benchmark, we consider $\alpha_{\rm WIMP}=0.0323$ and $m_{\chi}=700$ GeV. In this case, we obtain $\Omega_{\rm FO}h^{2}=0.08$. The left panel in Fig. 1 shows the relic abundance of WIMP due to Hawking radiation from PBHs $\Omega_{\rm ev}h^{2}$ according to the initial PBH mass, $M_{\rm in}$, for initial PBH density $\beta=1\times 10^{-20}$. The blue solid curve indicates prediction without the memory burden effect. The dotted curves show the prediction with the memory burden effect for various $q$ values. Relation $\Omega_{\rm FO}h^{2}\gg\Omega_{\rm ev}h^{2}$ is satisfied for $q=0.5-5\times 10^{-6}$. As the relic density of WIMPs via Hawking radiation should vanish for $t_{\rm ev}^{k}<t_{\rm FO}$, a lower limit of $M_{\rm in}$ appears in each curve. These lower limits can indicate shifts to a lighter mass owing to the memory burden effect via relation $t_{\rm ev}^{k}\simeq 1/\Gamma_{\rm BH}^{k}\propto(qM_{\rm in})^{3+2k}$ from Eqs. (12) and (13).
Figure 1: Relic abundance of DM considering WIMPs and PBHs. Relic abundance of WIMPs due to Hawking radiation from PBHs $\Omega_{\rm ev}h^{2}$ (left) and total relic abundance of DM $\Omega_{\rm DM}h^{2}$ (right) according to initial PBH mass $M_{\rm in}$.
Figure 2: Constraints on initial PBH density $\beta$ considering WIMPs and PBHs. Top left: $\beta$ and $\beta_{\rm c}$ according to $M_{\rm in}$ for $q=5\times 10^{-6}$. Top right: $\beta$ according to $M_{\rm in}$ for various $q$ values. Bottom left: $\beta$ according to mass of WIMPs, $m_{\chi}$. Bottom right: $\beta$ according to coupling of WIMPs, $\alpha_{\rm WIMP}$.
The right panel in Fig. 1 shows the total relic abundance of DM, $\Omega_{\rm DM}h^{2}$, according to the initial PBH mass, $M_{\rm in}$. As for the left panel, the blue solid curve indicates prediction without the memory burden effect. The dotted black horizontal line indicates the observed relic abundance, $\Omega_{\rm DM}h^{2}=0.12$. The remaining dotted curves show the prediction with the memory burden effect for various $q$ values. The relic density of WIMPs via FO is fixed at $\Omega_{\rm FO}h^{2}=0.08$, and relation $\Omega_{\rm DM}h^{2}\simeq\Omega_{\rm FO}h^{2}+\Omega_{\rm PBH}h^{2}$ is satisfied. Hence, the trends in the dotted curves are reflected by the relic density of survived PBHs with the memory burden effect, $\Omega_{\rm PBH}h^{2}\propto q\beta/M_{\rm in}^{1/2}$. The upper limit of $M_{\rm in}$ per dotted curve is due to requirement $\beta<\beta_{\rm c}\propto 1/(q^{5/2+k}M_{\rm in}^{1+k})$ (see also the top-left panel in Fig. 2). The relic abundance of DM is highly sensitive to the memory burden effect. For example, a small memory burden effect of $q\lesssim\mathcal{O}(10^{-3})$ is only allowed for $M_{\rm in}\lesssim\mathcal{O}(10^{6})$ g, $\beta=1\times 10^{-20}$, and $\Omega_{\rm FO}h^{2}=0.08$.
Figure 2 show the constraint on the initial PBH density, $\beta$. The top-left panel shows $\beta$ and $\beta_{\rm c}$ according to $M_{\rm in}$ for $q=5\times 10^{-6}$ as an example. This panel shows the upper limit of $M_{\rm in}$ with $\beta<\beta_{\rm c}$ for $q=5\times 10^{-6}$. The top-right panel shows $\beta$ according to $M_{\rm in}$ for various $q$ values. The upper limit of $M_{\rm in}$ in each curve is due to requirement $\beta<\beta_{\rm c}$. The bottom-left panel shows $\beta$ according to the mass of WIMPs, $m_{\chi}$. The upper limit of mass $m_{\chi}$ is obtained at $\Omega_{\rm DM}h^{2}=\Omega_{\rm FO}h^{2}$. As we fix $\Omega_{\rm DM}h^{2}=0.12$ and $\alpha_{\rm WIMP}=0.0323$, the upper limit, $m_{\chi}=858$ GeV, is obtained using Eq. (30). The bottom-right panel shows $\beta$ according to the coupling of WIMPs, $\alpha_{\rm WIMP}$. Like for $\beta$ according to $m_{\chi}$, the lower limit of $\alpha_{\rm WIMP}$ is caused by the relation in Eq. (30). As a strong memory burden effect yields long-lived PBHs, the initial PBH density, $\beta$, for $\Omega_{\rm DM}h^{2}=0.12$ decreases with increasing $q$ in all panels, except for the top-left panel in Fig. 2.
3.2 FIMPs and PBHs
Figure 3: Relic abundance of DM and constraint on initial PBH density for FIMPs and PBHs. Only figures that differ from the case considering WIMPs are shown. Top: relic abundance of FIMPs owing to Hawking radiation from PBHs, $\Omega_{\rm ev}h^{2}$, according to initial PBH mass $M_{\rm in}$. Bottom left: $\beta$ according to mass of FIMPs, $m_{\chi}$. Bottom right: $\beta$ according to coupling of FIMPs, $\alpha_{\rm FIMP}$.
Next, we consider DM made of FIMPs and PBHs. We assume that the masses of bath particles are negligible compared with those of FIMPs and that the FI mechanism is described by a pair production process with constant matrix element $|\mathcal{M}(f_{1}f_{2}\rightarrow\chi\chi)|^{2}=\alpha_{\rm FIMP}^{2}$. For a simple model of FIMPs [50, 38], the relic abundance of FIMPs, $\chi$, via FI is given by [50, 38]
$\Omega_{\rm FI}h^{2}=9.7\times 10^{19}\alpha_{\rm FIMP}^{2}.$ (33)
If the FI mechanism alone allows to produce DM, the observed relic abundance of DM, $\Omega_{\rm DM}h^{2}=0.12$, is consistent with $\alpha_{\rm FIMP}=3.52\times 10^{-11}$ [50].
Given the feeble interactions of FIMPs, the FIMPs emitted by Hawking radiation can always contribute to the relic abundance of DM. The present-day DM density is given by
$\Omega_{\rm DM}h^{2}$ $=\Omega_{\rm FI}h^{2}+\Omega_{\rm PBH}h^{2}+\Omega_{\rm ev}h^{2}$
$=9.7\times 10^{19}\alpha_{\rm FIMP}^{2}+4.2\times 10^{26}q\beta\left(\frac{M_{\rm P}}{M_{\rm in}}\right)^{1/2}+2.3\times 10^{5}C\beta\frac{m_{\chi}}{\rm GeV},$ (34)
where the first–third terms are due to FI, survived PBHs, and PBH evaporation, respectively.
Like for the WIMP case, we consider $\Omega_{\rm FI}h^{2}\gg\Omega_{\rm ev}h^{2}$. As a benchmark, we consider $\alpha_{\rm FIIMP}=2.872\times 10^{-11}$, $m_{\chi}=700$ GeV ($\Omega_{\rm FI}h^{2}=0.08$), and $\beta=1\times 10^{-20}$. For this benchmark, we obtain the same prediction of $\Omega_{\rm DM}h^{2}$ with $M_{\rm in}$, as shown in the right panel of Fig. 1. In addition, the prediction of $\beta$ with $M_{\rm in}$ for this benchmark considering FIMPs becomes the same as that considering WIMPs in Fig. 2. Thus, we omit the plots of $\Omega_{\rm DM}h^{2}$ $M_{\rm in}$ and $\beta$ according to $M_{\rm in}$ for the FIMP case.
The top panel in Fig. 3 shows the relic abundance of FIMPs due to Hawking radiation from PBHs, $\Omega_{\rm ev}h^{2}$, according to initial PBH mass $M_{\rm in}$ for $\beta=1\times 10^{-20}$. $\Omega_{\rm FO}h^{2}\gg\Omega_{\rm ev}h^{2}$ is satisfied for $q=0.5-5\times 10^{-6}$. Unlike the WIMP case, $\Omega_{\rm ev}h^{2}$ can always contribute to the relic abundance with no lower limit of $M_{\rm in}$.
The two bottom panels in Fig. 3 show the constraints on initial PBH density $\beta$ considering FIMPs. The bottom-left panel shows $\beta$ according to the mass of FIMPs, $m_{\chi}$. Unlike the WIMP case, no strong constraint on the mass of FIMPs is observed at $\Omega_{\rm DM}h^{2}=\Omega_{\rm FO}h^{2}$. The bottom-right panel shows $\beta$ according to the coupling of FIMPs, $\alpha_{\rm FIMP}$. Unlike the WIMP case, the upper limit of $\alpha_{\rm FIMP}$ is caused by the relation in Eq. (33).
Like for the WIMP case, a strong memory burden effect yields long-lived PBHs, and initial PBH density $\beta$ for $\Omega_{\rm DM}h^{2}=0.12$ decreases with increasing $q$ in the two bottom panels of Fig. 3.
4 Summary
The lifetime of PBHs has recently been discovered to be extended by the memory burden effect. Hence, light PBHs may currently survive and become candidates for DM. We assume that DM is made of thermally produced WIMPs (or FIMPs), WIMPs (or FIMPs) produced via Hawking radiation of PBHs, and survived PBHs via the memory burden effect. If the effect of thermal production of WIMPs (or FIMPs) is much smaller than that of particle production of PBHs via Hawking radiation, we can omit the contribution of thermally produced WIMPs (or FIMPs), obtaining the case already studied in [23]. Thus, we focus on the case that the effect of thermal production of WIMPs (or FIMPs) is much larger than that of particle production of PBHs via Hawking radiation.
Results from numerical calculations show that the relic abundance of DM is highly sensitive to the memory burden effect. For example, a small memory burden effect, $q\lesssim\mathcal{O}(10^{-3})$, is only allowed for $M_{\rm in}\lesssim\mathcal{O}(10^{6})$ g, $\beta=1\times 10^{-20}$, and $\Omega_{\rm FO}h^{2}=0.08$ if DM is made of thermal WIMPs, WIMPs from PBHs, and survived PBHs. A similar result is obtained if we consider FIMPs instead of WIMPs as the DM components.
Although the results of this study can be inferred through qualitative analyses, we believe that performing a quantitative analysis to confirm the qualitative prediction is important for increasing the accuracy and significance of the findings.
Acknowledgement
The author, T. K., thanks Anish Ghoshal for useful discussions.