### v3

# WIMP/FIMP dark matter and primordial black holes with memory burden effect

Abstract
The lifetime of primordial black holes (PBHs), which formed in the early universe, can be extended by the memory burden effect. Light PBHs may exist today and be candidates for dark matter (DM). We assume that DM is made of thermally produced weakly interacting massive particles (WIMPs), WIMPs produced via the Hawking radiation of PBHs, and PBHs that survived Hawking evaporation via the memory burden effect. Feebly interacting massive particles (FIMPs) are alternatives to WIMPs. This paper shows that a small memory burden is preferable if the thermal production of WIMPs or FIMPs is much larger than the effect of PBH particle production via Hawking radiation. In addition, we show that the lower limit of the DM mass, called the warm DM (WDM) constraint, decreases with the memory burden effect of PBHs. Results suggest that the WDM constraint is more effective for FIMPs than for WIMPs.
{history}
1 Introduction
The existence and production mechanism of dark matter (DM) are a long-standing mystery in cosmology and particle physics [1]. The traditional and most studied thermal DM production mechanism is freeze out (FO) [2]. In this mechanism, DM particles, $\chi$, reached thermal equilibrium with thermal bath particles, $f$, in the early universe through annihilation and pair production, $\chi\chi\leftrightarrow\bar{f}f$. Subsequently, the temperature of the bath particles drops below the mass of the DM particles and $\Gamma\lesssim H$ holds, where $\Gamma$ is the reaction rate of annihilation, $\chi\chi\rightarrow\bar{f}f$, and $H$ is the Hubble parameter. Then, the comoving number density of the DM particles remains constant. A typical weak-scale cross section leads to an observed relic abundance of DM, so the FO scenario is usually discussed in the context of weakly interacting massive particles (WIMPs). Although WIMPs are a leading candidate explanation for DM, they cannot be detected in numerous experiments. Thus, feebly interacting massive particles (FIMPs) [3, 4] have been proposed based on an idea opposite that of WIMPs. FIMPs are DM particles that never reached thermal equilibrium with thermal bath particles because of very weak coupling between them. Assuming a negligible initial density of FIMPs, the feeble interaction of DM particles with thermal bath particles leads to FIMP production during the expansion of the universe. This DM production scheme, called the freeze-in (FI) mechanism, is applicable to scenarios such as decay ($f_{1}\rightarrow f_{2}\chi$), scattering ($f_{1}f_{2}\rightarrow f_{3}\chi$), and pair production ($f_{1}f_{2}\rightarrow\chi\chi$).
Primordial black holes (PBHs), which formed in the early universe, are also feasible candidates for explaining DM [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. PBHs emit particles through Hawking radiation [18], which reduces their mass. Therefore, PBHs have a finite lifetime. However, if this lifetime is longer than the age of the universe, then some PBHs may not completely evaporate and may exist today.
PBHs that formed with a mass of less than approximately $10^{15}$ g were believed not to have survived to the present day. However, recent findings suggest that the lifetime of PBHs can be extended by the memory burden effect [19, 20, 21, 22, 23, 24]. Quantum corrections for PBHs due to the memory burden effect are being studied extensively to explain the relic abundance of PBHs [25, 28, 26, 27], gravitational waves (GWs) [29, 30, 31, 32, 33, 34, 28, 35, 36, 37], axion DM [38], baryon asymmetry [41, 39, 40], cosmic neutrinos [42, 43, 44, 45], compact stars [46], and phenomena beyond those in the standard model (SM) of particle physics [47].
We assume that DM consists of WIMPs/FIMPs and survived PBHs. As Hawking radiation is induced by gravity, PBHs evaporate into all particle species. WIMPs/FIMPs are also produced by Hawking radiation. Thus, we consider the following DM components:
  • thermally produced WIMPs/FIMPs
  • WIMPs/FIMPs produced by the Hawking radiation of PBHs
  • survived PBHs
Similar studies neglect the memory burden effect [49, 48, 50, 51, 52, 53, 54, 55, 56, 57]. Considering the memory burden effect, researchers attempted to explain the relic abundance of DM via the Hawking radiation of PBHs and survived PBHs [25]. However, they did not account for DM related to thermally produced particles.
We simultaneously consider thermally produced WIMPs/FIMPs and the memory burden effect of PBHs to explain the relic abundance of DM. We can omit the contribution of thermally produced WIMPs/FIMPs to the relic abundance of DM if the effect of their thermal production is much smaller than that of particle production by PBHs via Hawking radiation. In this case, the results are the same as those previously reported [25].
According to Ref. [25], we discuss a case where the semiclassical phase of Hawking evaporation ends before big bang nucleosynthesis (BBN) and the PBHs evaporate completely after the present time due to the memory burden effect (the so-called early-matter-dominated era does not exist). To examine a setup different from those in previous studies, we focus on a case where the effect of the thermal production of WIMPs/FIMPs considerably exceeds that of PBH particle production via Hawking radiation.
The remainder of this paper is organized as follows. Section 2 provides a brief review of PBHs with the memory burden effect. In Section 3, we discuss the constraints on PBHs from BBN, cosmic microwave background radiation (CMB), GWs, and warm DM (WDM).We show that the lower limit of the DM mass (the WDM constraint, which is related to structure formation in the universe) decreases with the memory burden effect of PBHs. In Section 4, we assess the combined effects of PBHs and thermally produced WIMPs/FIMPs on the relic abundance of DM. Finally, Section 5 provides a summary of this study and its findings.
2 PBH with memory burden effect
2.1 Time evolution
Neglecting the memory burden effect, the time evolution of PBH mass $M_{\rm BH}$ is given by
$\frac{dM_{\rm BH}}{dt}=-\epsilon\frac{M_{\rm P}^{4}}{M_{\rm BH}^{2}},$ (1)
where $M_{\rm P}\simeq 2.4\times 10^{18}$ GeV is the reduced Planck mass and
$\epsilon=\frac{27}{4}\frac{\pi g_{\ast}(T_{\rm BH})}{480},$ (2)
with $g_{\ast}(T_{\rm BH})$ being the effective number of degrees of freedom. The Hawking temperature, $T_{\rm BH}$, is defined as
$T_{\rm BH}=\frac{M_{\rm P}^{2}}{M_{\rm BH}}.$ (3)
After integrating Eq. (1) from the time of PBH formation, $t_{\rm in}$, to time $t$, the PBH mass at time $t$ is calculated as
$M_{\rm BH}=M_{\rm in}\left[1-\Gamma_{\rm BH}^{0}(t-t_{\rm in})\right]^{1/3},$ (4)
where $M_{\rm in}$ is the initial PBH mass. The decay width associated with PBH evaporation is estimated to be
$\Gamma_{\rm BH}^{0}=\frac{3\epsilon M_{\rm P}^{4}}{M_{\rm in}^{3}}.$ (5)
Assuming PBH formation in a radiation-dominated era, PBH lifetime $\tau^{0}$ is obtained as
$\tau^{0}=t_{\rm ev}^{0}-t_{\rm in}=\frac{1}{\Gamma_{\rm BH}^{0}}\simeq 2.4\times 10^{-28}\left(\frac{M_{\rm in}}{1{\rm g}}\right)^{3}\ {\rm s},$ (6)
by solving $M_{\rm BH}(t_{\rm ev}^{0})=0$, where $t_{\rm ev}^{0}$ is the evaporation time.
Considering the memory burden effect, the time evolution of the PBH mass is modified to become [25]
$\frac{dM_{\rm BH}}{dt}=-\frac{\epsilon}{[S(M_{\rm BH})]^{k}}\frac{M_{\rm P}^{4}}{M_{\rm BH}^{2}},$ (7)
where
$S=\frac{1}{2}\left(\frac{M_{\rm BH}}{M_{\rm P}}\right)^{2}=\frac{1}{2}\left(\frac{M_{\rm P}}{T_{\rm BH}}\right)^{2}$ (8)
is the black hole entropy. The parameter $k$ characterizes the efficiency of the back reaction effect, with $k=0$ indicating Hawking radiation without the memory burden effect.
The semiclassical era is assumed valid until the PBH mass reaches
$M_{\rm BH}=qM_{\rm in},$ (9)
where $0<q<1$. From Eq. (4), we define the final time for the semiclassical era as
$t_{q}=t_{\rm ev}^{0}(1-q^{3}).$ (10)
After integrating Eq. (7) from time $t_{q}$ to $t$, we obtain
$M_{\rm BH}=qM_{\rm in}\left[1-\Gamma_{\rm BH}^{k}(t-t_{q})\right]^{1/(3+2k)},$ (11)
where
$\Gamma_{\rm BH}^{k}=2^{k}(3+2k)\epsilon M_{\rm P}\left(\frac{M_{P}}{qM_{\rm in}}\right)^{3+2k}$ (12)
at $k\neq 0$ and $0<q<1$. We estimate the evaporation time by solving $M_{\rm BH}(t_{\rm ev}^{k})=0$ as [21, 25, 41]
$t_{\rm ev}^{k}$ $=$ $t_{q}+\frac{1}{\Gamma_{\rm BH}^{k}}\simeq(1-q^{3})\frac{M_{\rm in}^{3}}{3\epsilon M_{\rm P}^{4}}$ (13)
$+\frac{1}{2^{k}(3+2k)\epsilon M_{\rm P}}\left(\frac{qM_{\rm in}}{M_{P}}\right)^{3+2k}.$
Approximation ($t_{\rm ev}^{k}\simeq 1/\Gamma_{\rm BH}^{k}$) is adequate in some cases.
2.2 Initial density
The initial density of a PBH is
$\beta=\frac{\rho_{\rm BH}^{\rm in}}{\rho_{\rm r}^{\rm in}},$ (14)
where $\rho_{\rm BH}^{\rm in}$ and $\rho_{\rm r}^{\rm in}=\pi^{2}g_{\ast}(T_{\rm in})T_{\rm in}^{4}/30$ are the initial energy densities of the PBH and radiation, respectively. We assume that the PBH formed during a radiation-dominated era in the universe. Hence, the radiation energy density varies as
$\rho_{\rm r}(a)=\rho_{\rm r}^{\rm in}\left(\frac{a_{\rm in}}{a}\right)^{4}.$ (15)
As a PBH behaves as matter, its energy density evolves as
$\rho_{\rm BH}(a)=\rho_{\rm BH}^{\rm in}\left(\frac{a_{\rm in}}{a}\right)^{3}\times\begin{cases}1&{\rm for}\ a<a_{\rm q}\\ q&{\rm for}\ a>a_{\rm q}\\ \end{cases},$ (16)
where $a$ is a scale factor and instant mass reduction is assumed as $M_{\rm in}\rightarrow qM_{\rm in}$ at $a=a_{\rm q}$. According to $\rho_{\rm r}=\rho_{\rm BH}$, the radiation–PBH equivalent may occur at $a_{\rm BH}=a_{\rm in}/(q\beta)$. A PBH-dominant era is possible if $a_{\rm BH}<a_{\rm ev}$, which corresponds to
$\beta>\frac{a_{\rm in}}{qa_{\rm ev}}=\beta_{\rm c},$ (17)
where
$\beta_{\rm c}=\frac{1}{q^{5/2+k}}\left(\frac{M_{\rm P}}{M_{\rm in}}\right)^{1+k}\sqrt{\frac{(3+2k)2^{k}\epsilon}{8\pi\gamma}},$ (18)
where $\gamma\simeq 0.2$. We assume that the initial PBH mass is related to the horizon size during PBH formation as
$M_{\rm in}=\frac{4}{3}\pi\gamma\rho_{\rm r}^{\rm in}H_{\rm in}^{-3}.$ (19)
In addition, we use relations $H=1/(2t)$, $a\propto t^{1/2}$, and $H_{\rm in}^{2}=\rho_{\rm r}^{\rm in}/(3M_{\rm P}^{2})$ for a radiation-dominated universe and set $t_{\rm ev}^{k}\simeq 1/\Gamma_{\rm BK}^{k}$ using Eq. (12).
As addressed in Ref. [25], $\beta<\beta_{\rm c}$ is a case of interest because we assume that PBHs will evaporate completely after the present time due to the memory burden effect and PBHs that have survived until the present day are DM candidates (no early-matter-dominated era is considered). Hereinafter, $\beta<\beta_{\rm c}$ should hold.
2.3 Survived PBH density
We assume that PBHs will evaporate completely after the present time due to the memory burden effect. Thus, the density of survived PBHs is [58]
$\Omega_{\rm PBH}h^{2}=1.6\times 10^{8}\frac{g_{0}}{g_{\rm q}}\frac{\rho_{\rm BH}(a_{\rm q})}{T^{3}(a_{\rm q})}\ {\rm GeV^{-1}}.$ (20)
Using the relation
$\frac{\rho_{\rm BH}(a_{\rm q})}{T^{3}(a_{\rm q})}=\frac{q\rho_{\rm BH}^{\rm in}}{T_{\rm in}^{3}}=\frac{q\beta\rho_{r}^{\rm in}}{T_{\rm in}^{3}}$ (21)
and Eq. (19), we obtain (Eq. (3.37) in [25])
$\Omega_{\rm PBH}h^{2}=4.2\times 10^{26}q\beta\left(\frac{M_{\rm P}}{M_{\rm in}}\right)^{1/2}.$ (22)
2.4 Particle production
The number of particle species, $i$, from PBH evaporation is controlled by the relation between the particle mass, $m_{i}$, and the PBH formation temperature, which is
$T_{\rm in}=T_{\rm BH}(t_{\rm in})=\frac{M_{\rm P}^{2}}{M_{\rm in}}.$ (23)
If $m_{i}<T_{\rm in}$, then particle emission from a PBH occurs from $t_{\rm in}$ to $t_{\rm ev}$. If $m_{i}>T_{\rm in}$, then particle emission starts from time $t_{i}$, where $m_{i}=T_{\rm BH}(t_{i})$. The relic abundance of emitted particle DM is given by Eq. (3.33) in [25].
$\Omega_{\rm ev}h^{2}=1.8\times 10^{8}N_{\chi}\beta\left(\frac{M_{\rm P}}{M_{\rm in}}\right)^{3/2}\frac{m_{\chi}}{\rm GeV},$ (24)
where
$N_{\chi}=\frac{15\xi g_{\chi}\zeta(3)}{g_{\ast}(T_{\rm BH})\pi^{4}}\times\begin{cases}(1-q^{2})\left(\frac{M_{\rm in}}{M_{\rm P}}\right)^{2},&{\rm for}\ m_{\chi}<T_{\rm in}\\ \left(\frac{M_{\rm P}}{m_{\chi}}\right)^{2}-\left(\frac{qM_{\rm in}}{M_{\rm P}}\right)^{2},&{\rm for}\ m_{\chi}>T_{\rm in}\\ \end{cases},$ (25)
where $\xi=1$ for bosons and $\xi=3/4$ for fermions.
3 Constraints
3.1 BBN
We focus on a case where the semiclassical phase of PBH evaporation ends before BBN. The initial mass of the PBH, which is partially evaporated until the final time for the semiclassical era, $t_{q}$, can be obtained from Eqs. (5), (6), and (10) as
$M_{\rm in,t_{q}}=1.61\times 10^{9}\left(\frac{1}{1-q^{3}}\right)^{1/3}\left(\frac{t_{q}}{1{\rm s}}\right)^{1/3}~{\rm g},$ (26)
where we assume $t_{\rm in}\ll t_{\rm ev}^{0}$.
In this study, we set the time of BBN as $t_{\rm BBN}=1$ s for simplicity. We obtain
$M_{\rm in,t_{\rm BBN}}=1.61\times 10^{9}\left(\frac{1}{1-q^{3}}\right)^{1/3}~{\rm g}$ (27)
at $t_{q}=t_{\rm BBN}$. If $M_{\rm in}>M_{\rm in,t_{\rm BBN}}$, then the semiclassical evaporation of PBHs happens after BBN. Because $1\leq(1-q^{3})^{-1/3}\leq 1.055$ at $0<q\leq 0.9$, $M_{\rm in,t_{\rm BBN}}$ is almost independent of $q$ in this study. We use $M_{\rm in}\leq M_{\rm in,t_{\rm BBN}}=1\times 10^{9}$ g as the BBN constraint, which is the same as that without the memory burden effect [7, 59, 60, 61, 49].
3.2 CMB
Constraints on inflation impose a lower limit on the PBH mass. Given that the Hubble parameter at a time is less than or equal to the Hubble parameter during inflation, we obtain the lower limit of the initial mass of the PBH as $M_{\rm in}\gtrsim 0.1$ g [49, 7]. Because this constraint comes from PBH generation, it is independent of the memory burden effect. Although this constraint depends on the details of gravitational collapse and the inflation scenario, we regard $M_{\rm in}\gtrsim 0.1$ g as the CMB constraint by assuming a standard slow-roll scenario for simplicity [7].
3.3 GWs
PBHs are potential sources of GWs [62, 63, 64, 65, 66, 67, 68, 69, 70, 72, 71, 73, 74, 75, 76, 77, 78, 79, 80]. PBH formation is accompanied by GWs, and primordial curvature perturbations can induce GWs in the final stage of inflation, immediately after horizon re-entry.
GWs can be induced by sources other than primordial fluctuations [81, 82, 83]. PBHs are randomly distributed in space according to approximate Poisson statistics immediately after formation. Although PBH gas behaves similarly to pressureless dust, its spatially inhomogeneous distribution causes isocurvature density fluctuations. At a sufficiently large initial fraction of PBHs in the early-matter-dominated (PBH-dominated) era ($\beta>\beta_{\rm c}$ case), the initial isocurvature fluctuations become curvature perturbations. These curvature perturbations then act as sources of secondary GWs [81, 82, 85, 83, 86, 87, 84, 88, 89]. These induced GWs can be detected via future GW observations, such as those obtained using Hanford–Livingstone–Virgo–KAGRA [90, 91, 92, 93], the Einstein Telescope [94, 95], Cosmic Explorer [96], Laser Interferometer Space Antenna [97], Deci-Hertz Interferometer Gravitational-Wave Observatory [98, 99], and Big Bang Observer [100, 101].
GWs induced by PBHs with the memory burden effect have been extensively studied recently [29, 30, 31, 32, 33, 34, 28, 35, 36, 37]. Findings suggest that the GW spectrum amplitude is enhanced by the longer early PBH-dominated phase caused by the memory burden effect, assuming the existence of the early-matter-dominated era ($\beta>\beta_{\rm c}$).
In this study, PBHs are assumed to evaporate completely after the present time due to the memory burden effect and PBHs that have survived until the present day are regarded as DM candidates ($\beta<\beta_{\rm c}$ should hold). We assume that the early-matter-dominated era did not exist and PBHs never dominated the energy density in the history of the universe. In addition, we focus on a case where the effect of the thermal production of WIMPs/FIMPs substantially exceeds that of particle production by PBHs. In this case, GWs from PBHs almost behave as if the memory burden effect is nonexistent, so a small GW amplitude is expected (see $q=0$ curve in Fig. 3 in Ref. [29] for example).
Figure 1: WDM constraint with memory burden effect of PBHs. Left panel: lower limit of WDM mass vs. initial mass of PBH. The left (right) gray region is excluded from the CMB (BBN) constraint. Right panel: $q$ dependence on lower limit of $m_{\chi}$.
3.4 WDM
DM should be cold enough to avoid erasing small-scale structures via free streaming. A constraint on WDM or noncold DM is obtained via the study of the Lyman-$\alpha$ forest [102, 103, 104, 105, 49, 106].
In cases with negligible DM interaction with any particles in the thermal bath, DM particles can only lose energy (or momentum) by redshifting. The present redshifted momentum of DM, $p(t_{0})$, is [48]
$p(t_{0})$ $=$ $\frac{a(t_{\rm ev})}{a(t_{0})}p(t_{\rm ev})$ (28)
$=$ $\frac{a(t_{\rm ev})}{a(t_{\rm eq})}\frac{a(t_{\rm eq})}{a(t_{0})}p(t_{\rm ev})$
$=$ $\frac{a(t_{\rm ev})}{a(t_{\rm eq})}\frac{\Omega_{\rm r}}{\Omega_{\rm m}}p(t_{\rm ev}),$
where $\Omega_{\rm r}$ and $\Omega_{\rm m}$ are the density parameters of radiation and matter, respectively. Using $\rho_{\rm r}\propto a^{-4}$, we obtain
$\frac{a(t_{\rm ev})}{a(t_{\rm eq})}=\left(\frac{\rho_{\rm r}(T_{\rm eq})}{\rho_{\rm r}(T_{\rm ev})}\right)^{1/4}.$ (29)
We can use the relations
$\rho_{\rm r}(T_{\rm eq})$ $=$ $\rho_{\rm c}\left(\frac{a(T_{0})}{a(T_{\rm eq})}\right)^{3}=\rho_{\rm c}a(T_{\rm eq})^{-3},$
$\rho_{\rm r}(T_{\rm ev})$ $=$ $3M_{\rm P}^{2}H^{2}=\frac{3M_{\rm P}^{2}}{(2t_{\rm ev})^{2}}\simeq\frac{3M_{\rm P}^{2}}{4\tau^{2}},$
$a(T_{\rm eq})$ $=$ $a(t_{\rm eq})=\frac{\Omega_{\rm r}}{\Omega_{\rm m}},$ (30)
and $p(t_{\rm ev})\simeq\langle E\rangle\simeq 6T_{\rm in}$ for light DM ($m_{\rm DM}\ll T_{\rm in}$), where $\rho_{\rm c}=3H_{0}^{2}/(8\pi G)=1.05h^{2}\times 10^{-5}\ {\rm GeV\ cm^{-3}}$ is the critical density of the universe and $\tau$ is the PBH lifetime.
Without the memory burden effect, we set $\tau=\tau^{0}$ under the assumption of $t_{\rm in}\ll t^{0}_{\rm ev}$. In this case, the momentum of a DM particle today is
$p(t_{0})=1.53\times 10^{-12}\left(\frac{M_{\rm in}}{M_{\rm P}}\right)^{1/2}~{\rm GeV}$ (31)
at $g_{*}(T_{\rm BH})=106.75$, $\Omega_{\rm r}=5.38\times 10^{-5}$, and $\Omega_{\rm m}=0.315$ [107]. Assuming the WDM is heavier than 3.5 keV [108], the upper limit of the DM velocity is approximately $v_{\rm DM}\lesssim 1.8\times 10^{-8}$ [109]. From $p(t_{0})=m_{\rm DM}v_{\rm DM}$, we obtain the following WDM constraint on the DM mass and the initial mass of the PBH:
$m_{\rm DM}\gtrsim 4.07\times 10^{-2}\left(\frac{M_{\rm in}}{1{\rm g}}\right)^{1/2}~{\rm GeV}.$ (32)
Considering the memory burden effect, we can utilize the final time for the semiclassical era, $t_{\rm q}$, as an effective lifetime of PBHs for particle production. In this case, we set $\tau=\tau^{0}(1-q^{3})$ instead of $\tau=\tau^{0}$ from Eq. (10). The WDM constraint with the memory burden effect is
$m_{\rm DM}\gtrsim 4.07\times 10^{-2}(1-q^{3})^{1/2}\left(\frac{M_{\rm in}}{1{\rm g}}\right)^{1/2}~{\rm GeV}.$ (33)
Thus, the allowed minimum DM mass is reduced by the memory burden effect.
Figure 1 shows the nature of Eq.(33). The left panel shows the lower limit of the WDM mass against the initial mass of the PBH. The left (right) gray region is excluded from the CMB (BBN) constraint. The blue solid line indicates the prediction without the memory burden effect, whereas the dotted lines show predictions with the memory burden effect at q=0.0050.9q=0.005-0.9. The right panel shows $q$ dependence on the lower limit of $m_{\chi}$. As expected, a large memory burden effect of the PBH reduces the WDM constraint.
4 WIMP/FIMP DM and PBHs
4.1 WIMPs and PBHs
First, we consider DM made of WIMPs and PBHs. The FO of WIMPs occurs at [2]
$x_{\rm FO}=\frac{m_{\chi}}{T_{\rm FO}}\simeq 20,$ (34)
where $T_{\rm FO}$ is the FO temperature. We assume that FO occurs during a radiation-dominated era. The Hubble parameter is related to the radiation temperature and time as $H=(\pi^{2}g_{\ast}/30)T^{2}/(3M_{\rm P})=1/(2t)$. Thus, the FO time is
$t_{\rm FO}=\frac{1}{2}\sqrt{\frac{90}{\pi^{2}g_{\ast}}}\frac{M_{\rm P}}{T_{\rm FO}^{2}}.$ (35)
The relic abundance of WIMPs due to FO, $\chi$, is [48]
$\Omega_{\rm FO}h^{2}=0.76\frac{x_{\rm FO}}{g_{\ast}^{1/2}}\frac{s_{0}}{M_{\rm P}\rho_{\rm c}\langle\sigma v\rangle}h^{2},$ (36)
where $\langle\sigma v\rangle$ is the thermally averaged cross section at $\chi\chi\rightarrow\bar{f}f$ and $s_{0}=2.89\times 10^{3}\ {\rm cm^{-3}}$ is the entropy density. Considering the so-called WIMP miracle, we parameterize
$\langle\sigma v\rangle=\frac{\alpha_{\rm WIMP}^{2}}{m_{\chi}^{2}},$ (37)
where $\alpha_{\rm WIMP}^{2}$ is the effective coupling at $\chi\chi\leftrightarrow\bar{f}f$. The relic abundance of WIMPs is
$\Omega_{\rm FO}h^{2}=1.7\times 10^{-10}\frac{1}{\alpha_{\rm WIMP}^{2}}\left(\frac{m_{\chi}}{\rm GeV}\right)^{2}.$ (38)
If the FO mechanism alone produces DM, then the observed relic abundance, $\Omega_{\rm DM}h^{2}=0.12$ [110], is consistent with $\alpha_{\rm WIMP}\simeq 0.026$ and $m_{\chi}\simeq 1$ TeV.
The potential impact of PBH evaporation without the memory burden effect on the production of DM particles via thermal processes has been deeply investigated. In the context of the FO or FI mechanism, the authors in Ref. [50] considered fermionic DM interacting with SM particles via a spin-1 mediator. They focused on a case where the mediator is heavier than the DM but has a smaller mass than the reheating temperature.
In the case of the FO mechanism, the authors confirmed that DM particles produced from PBH evaporation thermalize with SM particles if the PBHs evaporate before FO. Therefore, PBHs do not affect the FO mechanism dynamics or contribute to the final relic abundance. Conversely, when PBH evaporation occurs after FO, two scenarios are possible.
In the first scenario, PBHs may dominate the universe’s energy density before evaporating. With PBHs dominating the universe’s energy density, PBHs act as a DM source by continuously injecting entropy into the visible sector, possibly affecting thermal production. If FO occurs during or before a PBH-dominated era, then PBH evaporation can alter standard FO predictions during PBH domination. However, unless particles from PBHs interact efficiently with thermal DM remnants and destroy FO predictions, PBHs cannot simultaneously affect FO production dynamics and contribute significantly to the relic density if they evaporate after FO. This is due to the presence of the mediator, which provides additional annihilation channels for DM and causes resonant interactions between DM and SM particles. The interaction between the evaporated DM particles and the preexisting relic abundance or between the evaporated DM particles and SM particles must be efficient for PBHs to dominate the universe’s energy density before evaporating. In this case, the Boltzmann equation should be solved at the level of phase-space distributions because of the nontrivial phase-space distribution of the evaporation products [50].
In the second scenario, PBHs never dominate the energy density. In this case, the FO mechanism and PBH evaporation can both contribute without affecting each other.
In this study, we focus on a case where the effect of the thermal production of WIMPs/FIMPs considerably exceeds that of PBH particle production via Hawking radiation. In addition, we assume that no early-matter-dominated era exists and PBHs never dominate the energy density in the history of the universe (we focus on a scenario where $\beta<\beta_{\rm c}$ holds). Therefore, DM particles from PBH evaporation are expected not to interact with thermally produced WIMPs.
To ensure that thermalization does not occur, we evaluate the interaction rates between particles from PBHs and thermally produced particles and compare them with the Hubble scale, as done in Ref. [50]. However, the current study does not examine the details of this problem. We take the range of $q$ at $\Omega_{\rm FO}h^{2}\gg\Omega_{\rm ev}h^{2}$.
We assume that the DM particle is a Majorana fermion (e.g., $\xi=3/4$ and $g_{\chi}=1$). In addition, we omit the negligible contribution of $g_{\chi}=1$ to the effective number of degrees of freedom and use $g_{*}=106.75$. The present-day density of DM is
$\Omega_{\rm DM}h^{2}$ $=\begin{cases}\Omega_{\rm FO}h^{2}+\Omega_{\rm PBH}h^{2}&{\rm for}\ t_{\rm ev}^{k}<t_{\rm FO}\\ \Omega_{\rm FO}h^{2}+\Omega_{\rm PBH}h^{2}+\Omega_{\rm ev}h^{2}&{\rm for}\ t_{\rm FO}<t_{\rm ev}^{k}\\ \end{cases},$
$=1.7\times 10^{-10}\frac{1}{\alpha_{\rm WIMP}^{2}}\left(\frac{m_{\chi}}{\rm GeV}\right)^{2}$
$\quad+4.2\times 10^{26}q\beta\left(\frac{M_{\rm P}}{M_{\rm in}}\right)^{1/2}$
$\quad+\begin{cases}0&{\rm for}\ t_{\rm ev}^{k}<t_{\rm FO}\\ 2.3\times 10^{5}C\beta\frac{m_{\chi}}{\rm GeV}&{\rm for}\ t_{\rm FO}<t_{\rm ev}^{k}\\ \end{cases},$ (39)
where the first–third terms are due to FO, survived PBHs, and PBH evaporation. The coefficient $C$ is
[…]
where $T_{\rm in}$ is given by Eq. (23).
We study the case of $\Omega_{\rm FO}h^{2}\gg\Omega_{\rm ev}h^{2}$. In the following numerical calculations, we assume $k=1$ as the simplest case. We regard $\alpha_{\rm WIMP}=0.0323$ and $m_{\chi}=700$ GeV as the benchmark. In this case, we obtain $\Omega_{\rm FO}h^{2}=0.08$.
Figure 2: Relic abundance of DM considering WIMPs and PBHs. Left: relic abundance of WIMPs due to Hawking radiation from PBHs, $\Omega_{\rm ev}h^{2}$. Right: total relic abundance of DM, $\Omega_{\rm DM}h^{2}$, according to initial PBH mass, $M_{\rm in}$. The left (right) gray region is excluded from the CMB (BBN) constraint.
The left panel in Fig. 2 shows the relic abundance of WIMPs due to Hawking radiation from PBHs, $\Omega_{\rm ev}h^{2}$, according to the initial PBH mass, $M_{\rm in}$, at the initial PBH density, $\beta=1\times 10^{-20}$. The left (right) gray region is excluded from the CMB (BBN) constraint. The blue solid line indicates the prediction without the memory burden effect, and the dotted lines show the predictions with the memory burden effect at q=0.0050.9q=0.005-0.9. $\Omega_{\rm FO}h^{2}\gg\Omega_{\rm ev}h^{2}$ ($\Omega_{\rm ev}/\Omega_{\rm FO}\lesssim 10^{-4}$) holds. As the relic density of WIMPs due to Hawking radiation should vanish at $t_{\rm ev}^{k}<t_{\rm FO}$, each curve has a lower limit of $M_{\rm in}$. These lower limits can indicate shifts to a lighter mass owing to the memory burden effect via $t_{\rm ev}^{k}\simeq 1/\Gamma_{\rm BH}^{k}\propto(qM_{\rm in})^{3+2k}$ from Eqs. (12) and (13).
The right panel in Fig. 2 shows the total relic abundance of DM, $\Omega_{\rm DM}h^{2}$, according to the initial PBH mass, $M_{\rm in}$. The blue solid line indicates the prediction without the memory burden effect, and the black horizontal dotted line indicates the observed relic abundance, $\Omega_{\rm DM}h^{2}=0.12$. The other dotted lines show the predictions with the memory burden effect at various $q$. The relic density of WIMPs due to FO is fixed at $\Omega_{\rm FO}h^{2}=0.08$, and $\Omega_{\rm DM}h^{2}\simeq\Omega_{\rm FO}h^{2}+\Omega_{\rm PBH}h^{2}$ holds. Hence, the trends of the dotted-line curves are reflected in the relic density of the survived PBHs with the memory burden effect, $\Omega_{\rm PBH}h^{2}\propto q\beta/M_{\rm in}^{1/2}$. The DM relic abundance is highly sensitive to the memory burden effect. For example, a small memory burden effect of $q\lesssim\mathcal{O}(10^{-3})$ is only allowed at $M_{\rm in}\lesssim\mathcal{O}(10^{6})$ g, $\beta=1\times 10^{-20}$, and $\Omega_{\rm FO}h^{2}=0.08$.
Figure 3 shows the constraint on the initial PBH density, $\beta$. The top panel shows $\beta$ against $M_{\rm in}$ at various $q$. The left (right) gray region is excluded from the CMB (BBN) constraint. The upper limit of $M_{\rm in}$ in each curve is due to the required $\beta<\beta_{\rm c}\propto 1/(q^{5/2+k}M_{\rm in}^{1+k})$.
The bottom-left panel shows $\beta$ against $m_{\chi}$. As a strong memory burden effect yields long-lived PBHs, the initial PBH density, $\beta$, decreases with an increase in $q$. The upper limits of $m_{\chi}$ in the curves are obtained at $\Omega_{\rm DM}h^{2}=\Omega_{\rm FO}h^{2}$. As we fix $\Omega_{\rm DM}h^{2}=0.12$ and $\alpha_{\rm WIMP}=0.0323$, the upper limit, $m_{\chi}=858$ GeV, is obtained using Eq. (38). The lower limits of $m_{\chi}$ in the curves are controlled by the required $\beta<\beta_{\rm c}$. We obtain
$\beta\simeq\frac{\sqrt{M_{\rm in}}}{4.2\times 10^{26}q\sqrt{M_{\rm P}}}\left(\Omega_{\rm DM}h^{2}-1.7\times 10^{-10}\frac{m_{\chi}^{2}}{\alpha_{\rm WIMP}^{2}}\right),$ (41)
at $\Omega_{\rm FO}h^{2}\gg\Omega_{\rm ev}h^{2}$ from Eq.(39). Thus, a small $m_{\chi}$ with a large $q$ increases $\beta$ and eliminates it via $\beta<\beta_{\rm c}$. In principle, the WDM constraint may also contribute to the lower limit of $m_{\chi}$. However, the WDM constraint is weaker than the $\beta<\beta_{\rm c}$ constraint.
The bottom-right panel shows $\beta$ according to the coupling of WIMPs, $\alpha_{\rm WIMP}$. The lower limit of $\alpha_{\rm WIMP}$ is caused by Eq. (38). The upper limits of $\alpha_{\rm WIMP}$ in the curves are obtained from the $\beta<\beta_{\rm c}$ constraint. Because $\beta\propto m_{\chi}^{2}/\alpha_{\rm WIMP}^{2}$, the bottom-left ($\beta$ vs. $m_{\chi}$) and bottom-right ($\beta$ vs. $\alpha_{\rm WIMP}$) panels are essentially reversed.
Figure 3: Constraints on initial PBH density, $\beta$, considering WIMPs and PBHs. Top: $\beta$ vs. $M_{\rm in}$ at various $q$. The left (right) gray region is excluded from the CMB (BBN) constraint. Bottom left: $\beta$ v.s, $m_{\chi}$. Bottom right: $\beta$ vs. $\alpha_{\rm WIMP}$.
Figure 4: Relic abundance of DM and constraint on initial PBH density for FIMPs and PBHs. Only figures that differ from the case considering WIMPs are shown. Top: relic abundance of FIMPs owing to Hawking radiation from PBHs, $\Omega_{\rm ev}h^{2}$, according to initial PBH mass, $M_{\rm in}$. The left (right) gray region is excluded from the CMB (BBN) constraint. Bottom left: $\beta$ according to FIMP mass, $m_{\chi}$. Bottom right: $\beta$ according to FIMP coupling, $\alpha_{\rm FIMP}$.
4.2 FIMPs and PBHs
Next, we consider DM consisting of FIMPs and PBHs. We assume that the masses of thermal bath particles are negligible compared with those of FIMPs and that the FI mechanism is described by pair production with a constant matrix element, $|\mathcal{M}(f_{1}f_{2}\rightarrow\chi\chi)|^{2}=\alpha_{\rm FIMP}^{2}$. For a simple model of FIMPs, the relic abundance of FIMPs, $\chi$, due to FI is [111, 48]
$\Omega_{\rm FI}h^{2}=9.7\times 10^{19}\alpha_{\rm FIMP}^{2}.$ (42)
If the FI mechanism alone produces DM, then the observed relic abundance of DM, $\Omega_{\rm DM}h^{2}=0.12$, is consistent with $\alpha_{\rm FIMP}=3.52\times 10^{-11}$ [111].
According to Ref. [50], in the case of the FI mechanism, particle interactions are small enough to avoid thermalization between particles from PBHs (FIMPs) and SM particles. If PBHs dominate the universe’s energy density before evaporating, then the annihilation cross section should be several orders of magnitude larger than that in the standard FI case to ensure efficient thermalization.
In this study, PBHs never dominate the energy density in the history of the universe. DM particles from PBH evaporation are expected not to interact with thermally produced FIMPs, as in the WIMP case. To ensure that thermalization does not occur in the FI case, we compare the interaction rates between particles from PBHs and thermally produced particles with the Hubble scale [50]. However, instead of this estimation, we consider the range of parameters $q$ at $\Omega_{\rm FI}h^{2}\gg\Omega_{\rm ev}h^{2}$.
Given the feeble interactions of FIMPs, the FIMPs emitted by Hawking radiation can always contribute to the relic abundance of DM. The present-day DM density is
$\Omega_{\rm DM}h^{2}$ $=\Omega_{\rm FI}h^{2}+\Omega_{\rm PBH}h^{2}+\Omega_{\rm ev}h^{2}$
$=9.7\times 10^{19}\alpha_{\rm FIMP}^{2}+4.2\times 10^{26}q\beta\left(\frac{M_{\rm P}}{M_{\rm in}}\right)^{1/2}$
$+2.3\times 10^{5}C\beta\frac{m_{\chi}}{\rm GeV},$ (43)
where the first–third terms are due to FI, survived PBHs, and PBH evaporation.
As in the WIMP case, we consider $\Omega_{\rm FI}h^{2}\gg\Omega_{\rm ev}h^{2}$. We regard $\alpha_{\rm FIIMP}=2.872\times 10^{-11}$, $m_{\chi}=700$ GeV ($\Omega_{\rm FI}h^{2}=0.08$), and $\beta=1\times 10^{-20}$ as the benchmark. For this benchmark, we obtain the same prediction of $\Omega_{\rm DM}h^{2}$ with $M_{\rm in}$ (right panel of Fig. 2). Thus, we omit the plot of $\Omega_{\rm DM}h^{2}$ - $M_{\rm in}$ for the FIMP case.
The top panel in Fig. 4 shows the relic abundance of FIMPs due to Hawking radiation from PBHs, $\Omega_{\rm ev}h^{2}$, according to the initial PBH mass, $M_{\rm in}$, at $\beta=1\times 10^{-20}$ and q=0.0050.9q=0.005-0.9. $\Omega_{\rm FI}h^{2}\gg\Omega_{\rm ev}h^{2}$ ($\Omega_{\rm ev}/\Omega_{\rm FI}\lesssim 10^{-4}$) holds. The left (right) gray region is excluded from the CMB (BBN) constraint.
The two bottom panels in Fig. 4 show the constraints on the initial PBH density, $\beta$, considering FIMPs. As in the WIMP case, a strong memory burden effect yields long-lived PBHs, and the initial PBH density, $\beta$, decreases with an increase in $q$ in the two bottom panels of Fig. 4.
The bottom-left panel shows $\beta$ according to the FIMP mass, $m_{\chi}$. Unlike in the WIMP case, the lower limits of $m_{\chi}$ in the curves are controlled by the WDM constraint (these lower limits are controlled by the $\beta<\beta_{\rm c}$ constraint in the WIMP case). From Eq.(43), we obtain
$\beta\simeq\frac{\sqrt{M_{\rm in}}}{4.2\times 10^{26}q\sqrt{M_{\rm P}}}\left(\Omega_{\rm DM}h^{2}-9.7\times 10^{19}\alpha_{\rm FIMP}^{2}\right),$ (44)
at $\Omega_{\rm FI}h^{2}\gg\Omega_{\rm ev}h^{2}$ in the FIMP case. Therefore, $\beta$ is independent of $m_{\chi}$ and the relation $\beta<\beta_{\rm c}$ does not constrain $m_{\chi}$. We obtain $21.2\leq m_{\chi}^{\rm min}[{\rm GeV}]\leq 40.7$ at $0.9\geq q\geq 0$ with $M_{\rm in}=1\times 10^{6}$ from the WDM constraint.
The bottom-right panel shows $\beta$ according to FIMP coupling, $\alpha_{\rm FIMP}$. Unlike in the WIMP case, the upper limit of $\alpha_{\rm FIMP}$ is caused by Eq. (42). The lower limit of $\alpha_{\rm FIMP}$ is obtained through the $\beta<\beta_{\rm c}$ condition. A small $\alpha_{\rm FIMP}$ with a large $q$ increases $\beta$ and eliminates it via the $\beta<\beta_{\rm c}$ condition.
5 Summary
The lifetime of PBHs is extended by the memory burden effect. Hence, light PBHs may survive and become candidates for DM. We assume that DM is made of thermally produced WIMPs/FIMPs, WIMPs/FIMPs produced via the Hawking radiation of PBHs, and PBHs that survived via the memory burden effect. We can omit the contribution of thermally produced WIMPs/FIMPs to DM relic abundance, obtaining the case studied by [25], if the effect of the thermal production of WIMPs/FIMPs is much smaller than that of PBH particle production via Hawking radiation. Thus, we focus on a case where the effect of the thermal production of WIMPs/FIMPs substantially exceeds that of PBH particle production via Hawking radiation.
The relic abundance of DM is highly sensitive to the memory burden effect. For example, a small memory burden effect, $q\lesssim\mathcal{O}(10^{-3})$, is only allowed at $M_{\rm in}\lesssim\mathcal{O}(10^{6})$ g, $\beta=1\times 10^{-20}$, and $\Omega_{\rm FO}h^{2}=0.08$ for DM consisting of thermal WIMPs, WIMPs from PBHs, and survived PBHs. A similar result is obtained if we consider FIMPs instead of WIMPs as the DM components.
We also find that the WDM constraint–the lower limit of the DM mass, which is related to structure formation in the universe–decreases with the memory burden effect of PBHs. The WDM constraint is more effective for FIMPs than for WIMPs.
Acknowledgement
The author, T. K., thanks Anish Ghoshal for useful discussions.